OpenStudy (anonymous):

How do I solve this problem? (3(2))^2

OpenStudy (anonymous):

try 36.

OpenStudy (anonymous):

3*2=6 and the equation is squared, so 6*6=36

OpenStudy (anonymous):

3*2=6 6*6=36